OpenStudy (anonymous):

Find all solutions to the equation in the interval [0,2pi): sin(x+pi)-sin(x)-1=0

OpenStudy (anonymous):

sin(x+pi)-sinx+1=0 sinx cos pi + cosx sin pi - sinx +1 =0 -sinx-sinx+1=0 -2sinx=-1 sinx=1/2 x=pi/6+2npi or x=5pi/6+2npi P.S ,cos pi =-1 ,and sin pi =0 cos(x+pi)-cosx-1=0 cosx cos pi -sinx sin pi - cosx-1=0 -cosx-cosx-1=0 -2cosx=1 cosx=-1/2 x=-pi/3+npi sin(x+pi/6)-sin(x-pi/6)=1/2 sinx sin pi/6 + cosx cos pi/6 - (sinx sin pi/6 - cosx cos pi/6)=1/2 sinx sin pi/6 + sqrt3 /2 cosx - sinx sin pi/6 + sqrt3 /2 cosx =1/2 sqrt3cosx=1 cosx=1/2sqrt3 x=73 degrees hope this helps :)

OpenStudy (anonymous):

thank you!!